# Mining REST APIs for Potential Mass Assignment Vulnerabilities

### v1

@@ Abstract.
REST APIs have a pivotal role in accessing protected resources within cyberspace. Despite the availability of security testing tools, mass assignment vulnerabilities are common, yielding unauthorized access to sensitive data. We propose a lightweight approach to mine the REST API specifications and identify operations and attributes that are prone to mass assignment. We conducted a preliminary study on 100 APIs and found 25 prone to this vulnerability. We confirmed nine real vulnerable operations in six open-source APIs.
@@
journalyear: 2024conference: 28th International Conference on Evaluation and Assessment in Software Engineering; June 18–21, 2024; Salerno, Italybooktitle: 28th International Conference on Evaluation and Assessment in Software Engineering (EASE 2024), June 18–21, 2024, Salerno, Italycopyright: nonedoi: 10.1145/xxxccs: Security and privacy Web application security
@@ 1. Introduction
Mass assignment is a critical but overlooked vulnerability in REST APIs. It occurs when REST APIs allow the unintended modification of attributes, often yielding unauthorized access to sensitive data. This vulnerability arises due to an incorrect configuration of widely used REST API frameworks that typically facilitate automatic binding between input data fields and the internal data representation, such as database columns.
The support for identifying mass assignment vulnerabilities in REST APIs is limited. Akto (Team, 2023a) and RestTestGen (Team, 2023b) are two examples of tools for detecting mass assignment vulnerabilities in REST APIs. RestTestGen is an automated black-box testing tool, and Akto is semi-automated. Nonetheless, existing tools evaluate a running API, and developers have no support to uncover mass assignment vulnerabilities in earlier development stages.
@@ 2.1. REST APIs and OpenAPI specifications
Every REST API should follow the OpenAPI standard to describe the API’s structure and behavior. In particular, REST APIs include an OpenAPI specification file, typically structured in JSON or YAML format, which describes the endpoints, operations and their attributes, as well as the request and response schemas.
Listing 1 presents a snippet of the OpenAPI specification for a “task management” API. Following an initial header specifying versions, licenses, and the API’s base URL, this specification features an array of paths representing the available URI endpoints in the API. In this example, the HTTP URI leading to a task resource is /tasks (line 12), and the HTTP operations GET and POST (lines 13 and 28) are utilized to retrieve the list of existing tasks and create a new task in the system, respectively. These operations have common attributes such as title (lines 22 and 35) and assignee (lines 24 and 37), which delineate the task’s title and the person responsible for it.
@@ 2.2. Mass Assignment Vulnerability
Developers usually rely on frameworks to build REST APIs. These frameworks, such as Spring for Java, Flask for Python, Express.js for JavaScript, and Laravel for PHP, offer a suite of reusable components and features to facilitate REST API development. One of the features typically provided by these frameworks is called auto-binding, a mechanism that adopts naming conventions to automatically map input data in HTTP requests (i.e., attributes) to the backend data objects (e.g., database columns) when they share the same name. This feature is typically enabled by default for all attributes. A mass assignment vulnerability, also known as “object injection” or “auto-binding vulnerability”, occurs when developers neglect to disable this feature for attributes that are meant to be read-only. Therefore, an attacker can add an extra attribute to an HTTP request (one that was not intended to be changed), and the auto-binding feature would automatically link that attribute to its corresponding database column. In principle, this attribute is neither part of the API specification nor the API documentation, and it should have not been processed. Nevertheless, the attacker who exploits this feature, will be able to manipulate and alter data in the database, posing a significant security risk.
For instance, consider the specification of the “Task Management” API shown in Listing 1. Suppose the JSON task object within an HTTP request maps to the tasks table in the database, which includes a (supposedly) read-only boolean column named status and two modifiable columns, namely title and assignee. If the REST framework lacks a proper configuration, it may automatically link an additional status attribute in a “create task” request to the corresponding status column in the tasks table. That is, an attacker could manipulate a request body of the POST /tasks operation by introducing an extra status attribute not specified in the OpenAPI specification. The framework would then automatically associate this additional attribute with the status column in the tasks table, allowing the attacker to overwrite the legitimate value in the database with the manipulated HTTP attribute value.
@@ 22 Return MassList
LightMass identifies similar operations based on similar attributes among operations. Firstly, to facilitate a uniform comparison, Porter’s stemming algorithm (Porter, 1980) is employed to standardize attribute names, i.e., reducing attribute names to their core or root forms. Secondly, it utilizes the Jaccard coefficient to identify similar operations:
@@ 5. Related Work
Gadient et al. (Gadient et al., 2020) mined 9,714 Web APIs from 3,376 mobile apps, and found that in 500 apps, these APIs transmit embedded code (e.g., SQL and JavaScrip commands), exposing the app users and web servers to code injection attacks. In a follow-up study (Gadient et al., 2021), they also discovered that API servers are usually misconfigured. They observed that on average every second server suffers from version information leaks, and worryingly, servers are usually set up once and never touched again, yielding severe security risks.
@@
The number of studies on mass assignment vulnerability is limited. Corradini et al. (Corradini et al., 2023) developed an automated black-box testing approach to find mass assignment vulnerabilities in RESTful APIs. It uses EM clustering to group operations within the endpoints. Subsequently, abstract testing templates are instantiated to automatically generate interaction sequences, to exploit potential vulnerabilities. This tool requires the API to be in a running status for interaction with HTTP operations.

### v2

# Preprint Version

@@ Abstract.
REST APIs have a pivotal role in accessing protected resources. Despite the availability of security testing tools, mass assignment vulnerabilities are common in REST APIs, leading to unauthorized manipulation of sensitive data. We propose a lightweight approach to mine the REST API specifications and identify operations and attributes that are prone to mass assignment. We conducted a preliminary study on 100 APIs and found 25 prone to this vulnerability. We confirmed nine real vulnerable operations in six APIs.
@@
journalyear: 2024copyright: rightsretainedconference: 28th International Conference on Evaluation and Assessment in Software Engineering; June 18–21, 2024; Salerno, Italybooktitle: 28th International Conference on Evaluation and Assessment in Software Engineering (EASE 2024), June 18–21, 2024, Salerno, Italydoi: 10.1145/3661167.3661204isbn: 979-8-4007-1701-7/24/06copyright: nonedoi: 10.1145/xxxccs: Security and privacy Web application security
@@ 1. Introduction
Mass assignment is a critical but overlooked vulnerability in REST APIs. It occurs when REST APIs allow the unintended modification of attributes, leading to unauthorized manipulation of sensitive data. This vulnerability arises due to an incorrect configuration of widely used REST API frameworks that typically facilitate automatic binding between input data fields and the internal data representation, such as database columns.
The support for identifying mass assignment vulnerabilities in REST APIs is limited. Akto (Team, 2023a) and RestTestGen (Team, 2023b) are two examples of tools for detecting mass assignment vulnerabilities in REST APIs. RestTestGen is an automated black-box testing tool, and Akto is semi-automated. Nonetheless, existing tools mostly evaluate a running API, support to uncover mass assignment vulnerabilities in earlier development stages is limited.
@@ 2.1. REST APIs and OpenAPI specifications
Developers usually follow the OpenAPI standard to describe the API’s structure and behavior. In particular, the OpenAPI specification file, typically structured in JSON or YAML format, describes the endpoints, operations and their attributes, as well as the request and response schemas.
Listing 1 presents a snippet of the OpenAPI specification for a Task Management API. Following an initial header specifying versions, licenses, and the API’s base URL, this specification features an array of paths representing the available URI endpoints in the API. In this example, the HTTP URI leading to a task resource is /tasks (line 12), and the HTTP operations GET and POST (lines 13 and 28) are utilized to retrieve the list of existing tasks and create a new task in the system, respectively. These operations have common attributes such as title (lines 22 and 35) and assignee (lines 24 and 37), which delineate the task’s title and the person responsible for it.
@@ 2.2. Mass Assignment Vulnerability
Developers usually rely on frameworks to build REST APIs. These frameworks, such as Spring for Java, Flask for Python, Express.js for JavaScript, and Laravel for PHP, offer a suite of reusable components and features to facilitate REST API development. One of the features typically provided by these frameworks is called auto-binding, a mechanism that adopts naming conventions to automatically map input data in HTTP requests (i.e., parameters) to the backend data objects (e.g., database columns) when they share the same name. This feature is typically enabled by default for all attributes. A mass assignment vulnerability, also known as “object injection” or “auto-binding vulnerability”, occurs when developers neglect to disable this feature for attributes that are meant to be read-only. Therefore, an attacker can add an extra attribute to an HTTP request (one that was not intended to be changed), and the auto-binding feature would automatically link that attribute to its corresponding internal representation, for example, a database column. In principle, this attribute is neither part of the API specification nor the API documentation, and it should have not been processed. Nevertheless, the attacker who exploits this feature, will be able to manipulate and alter data in the database, posing a significant security risk.
For instance, consider the specification of the Task Management API shown in Listing 1. Suppose the JSON task object within an HTTP request maps to the tasks table in the database, which includes a (supposedly) read-only boolean column named status and two modifiable columns, namely title and assignee. If the REST framework lacks a proper configuration, it may automatically link an additional status attribute in a “create task” request to the corresponding status column in the tasks table. That is, an attacker could manipulate a request body of the POST /tasks operation by introducing an extra status attribute not specified in the OpenAPI specification. The framework would then automatically associate this additional attribute with the status column in the tasks table, allowing the attacker to overwrite the legitimate value in the database with the manipulated HTTP attribute value.
@@ 22 Return MassList
LightMass identifies similar operations based on similar attributes. Firstly, to facilitate a uniform comparison, Porter’s stemming algorithm (Porter, 1980) is employed to standardize attribute names, i.e., reducing attribute names to their core or root forms. Secondly, it utilizes the Jaccard coefficient to identify similar operations:
@@ 5. Related Work
Gadient et al. (Gadient et al., 2020) mined 9,714 Web APIs from 3,376 mobile apps, and found that in 500 apps, these APIs transmit embedded code (e.g., SQL and JavaScrip commands), exposing the app users and web servers to code injection attacks. In a follow-up study (Gadient et al., 2021), they also discovered that API servers are usually misconfigured. They observed that on average every second server suffers from version information leaks, and worryingly, servers are typically set up once and then left untouched for up to fourteen months, yielding severe security risks.
@@
The number of studies on mass assignment vulnerability is limited. Corradini et al. (Corradini et al., 2023) developed an automated black-box testing approach to find mass assignment vulnerabilities in RESTful APIs. The approach, built on top of the RestTestGen framework (Corradini et al., 2022), uses EM clustering to group operations within the endpoints. Subsequently, abstract testing templates are instantiated to automatically generate interaction sequences, to exploit potential vulnerabilities. This tool requires the API to be in a running status for interaction with HTTP operations.